OpenStudy (aripotta):

check to see if this is right please? put in general form: y = x^2 + 14x + 29 y = x^2 + 14x + 29 y - 29 = x^2 + 14x y - 29 = x^2 + 14x + 49 y + 20 = (x + 7)^2 y = (x + 7)^2 - 20

pooja195 (pooja195):

y = x² + 14x + 29 y-29=x²+14x y-29+49=(x+7)² y=(x+7)² - 20

pooja195 (pooja195):

CORRRRRCET!!!! :)

OpenStudy (aripotta):

thanks :]

pooja195 (pooja195):

np!